OpenStudy (anonymous):

5^x=4 What are the exact and approximate solutions?

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

in your case \(b=5A=4\)

OpenStudy (anonymous):

ok i see and the aprox answer would be 0.8614 right